archive-es.com » ES » O » OVH.ES

Total: 788

Choose link from "Titles, links and description words view":

Or switch to "Titles and links view".
  • OVH Tasks  
    Paris Attached to Project ADSL SDSL Task Type Maintenance Category Backend Core Status CLOSED Percent Complete Details Maintenance is scheduled on our components in relation to task FS 13935 No customer impact expected Date Wednesday 22 July 2015 11 05AM

    Original URL path: http://status.ovh.es/?do=details&id=10071 (2015-10-15)
    Open archived version from archive


  • OVH Tasks  
    SDSL Task Type Maintenance Category Backend Core Status CLOSED Percent Complete Details Maintenance is scheduled on the internet network FS 13935 POP TH2 FR EU The LNS 1 PAR is connected to this network Packet loss may occur Date Wednesday

    Original URL path: http://status.ovh.es/?do=details&id=10070 (2015-10-15)
    Open archived version from archive

  • OVH Tasks  
    impacting the security of the loop of Marseille No customer impact We have opened a ticket for our fiber operator CCNs BEA13 CAV13 COR13 DRA13 FER13 GAR13 JOL13 MEN13 NAP13 PRA13 SMG13 SML13 STM13 Date Wednesday 22 July 2015 11 04AM Reason for closing Done Comments 2 Comment by OVH Wednesday 22 July 2015 11 04AM The fibre operator found the cut in an urban zone The intervention can only

    Original URL path: http://status.ovh.es/?do=details&id=10035 (2015-10-15)
    Open archived version from archive

  • OVH Tasks  
    Core Status CLOSED Percent Complete Details There s a general electrical issue on all NRA PPE75 Pumps Date Wednesday 22 July 2015 11 03AM Reason for closing Done Comments 1 Comment by OVH Friday 03 July 2015 10 43AM NRAs

    Original URL path: http://status.ovh.es/?do=details&id=10024 (2015-10-15)
    Open archived version from archive

  • OVH Tasks  
    Details A fibre outage impacted the loop in Paris We ve opened a ticket with our fibre operator Date Friday 12 June 2015 12 27PM Reason for closing Done Comments 2 Comment by OVH Tuesday 09 June 2015 13 20PM The fibre outage caused the network to behave abnormally It s been fixed We will continue to monitor the network until the optical loop is secure Comment by OVH Friday

    Original URL path: http://status.ovh.es/?do=details&id=9714 (2015-10-15)
    Open archived version from archive

  • OVH Tasks  
    Percent Complete Details DSLAM BNE75 BOB75 DAN75 GOB75 LIT75 RAS75 are currently inaccessible because of planned work being carried out by SFR on two points of the optcial loop The work should finish before 6am GMT 1 Date Tuesday 09

    Original URL path: http://status.ovh.es/?do=details&id=9712 (2015-10-15)
    Open archived version from archive

  • OVH Tasks  
    Category Backend Core Status CLOSED Percent Complete Details We ve lost contact with this component We re investigating Date Friday 12 June 2015 12 25PM Reason for closing Done Comments 1 Comment by OVH Friday 12 June 2015 12 25PM

    Original URL path: http://status.ovh.es/?do=details&id=9705 (2015-10-15)
    Open archived version from archive

  • OVH Tasks  
    Category Backend Core Status CLOSED Percent Complete Details SFR eligibility web services are currently experiencing issues ovhtelecom fr eligibility and orders are impacted Date Friday 29 May 2015 15 02PM Reason for closing Done Comments 1 Comment by OVH Friday

    Original URL path: http://status.ovh.es/?do=details&id=9622 (2015-10-15)
    Open archived version from archive



  •